# Degenerate scalar scenario of two Higgs doublet model with a complex singlet scalar

### v1

@@ III.2 Type-II, Type-Y
Therefore, in the degenerate limit of the mediator masses, the DM-quark scattering is suppressed as long as the potential (1) is used. In other words, the mixing term between $\Phi_{2}$ and $S$ is constrained by the degenerate scalar scenario. If mixing terms such as $\Phi_{2}^{\dagger}\Phi_{2}S,\Phi_{2}^{\dagger}\Phi_{2}S^{2}$ are added, the suppression mechanism does not work [^2] A more detailed analysis, including one-loop scattering, is given in Ref. Cho and Idegawa (2023). This discussion can be applied almost directly to the case of the 2HDMS..
Before finishing this section, we mention the usefulness of the degenerate scalar scenario in the Higgs search experiments at the LHC. In the 2HDMS, the couplings between the Higgs boson $H_{i}$ and the SM particle $X$ are given by the SM couplings multiplied by the mixing matrix elements $O_{1i}$. Taking the decay widths of $H_{i}$ as an example, we obtain
@@
where $\Gamma_{h\rightarrow XX}^{\mathrm{SM}}\left(m_{H_{i}}\right)$ is the decay width of the SM Higgs $h$ as a function of $m_{H_{i}}$. When we add up all the decay processes,
@@ IV Qualitative discussion for electroweak phase transition
We discuss the compatibility of the degenerate scalar scenario and EWPT qualitatively. As mentioned in Sec. III, the orthogonality of the mixing matrix (32) is important for the degenerate scalar scenario. Here, using the elements of the mixing matrix, the mixing term for SU(2)L doublet-singlet can be expressed as
@@
Compared to Eqs. (50) and (53), it can be seen that the DM-quark scattering is controlled by $\delta_{1}$ and $\delta_{2}$. In other words, the degenerate scalar scenario is realized by $\delta_{1}$ and $\delta_{2}$.
Strong first-order EWPT is essential for EWBG, and the condition is Arnold and McLerran (1987); Bochkarev and Shaposhnikov (1987); Funakubo and Senaha (2009);
@@
where $T_{C}$ is the critical temperature at which the effective potential has two degenerate minima, and $v_{C}$ is the nonzero Higgs VEV at $T=T_{C}$. Possible origins of first-order EWPT are the finite-temperature boson loop and the structure of the tree-level potential. In the former case, the addition of multiple bosons is effective, and the 2HDM specializes in this type of EWPT. On the other hand, in the latter case, the existence of a single scalar is important, and the CxSM causes this type of EWPT. Since the 2HDMS has multiple bosons and a singlet scalar that does not exist in the SM, EWPT of both types can occur. First, we qualitatively analyze tree-level driven EWPT using the so-called high temperature (HT) potential, which is composed of the tree-level potential and thermal masses.
EWPT pattern of 2HDMS is $\quantity(\langle\Phi_{1}\rangle,\langle\Phi_{2}\rangle,\langle S\rangle)=(0,0,v_{S}^{\prime})\to(v_{1},v_{2},v_{S})$, where $\langle\Phi_{i}\rangle$ and $\langle S\rangle$ are defined as the VEVs of the doublet fields $\Phi_{i}$ and the singlet field $S$, respectively. To discuss EWPT, we take the Landau gauge and denote the classical background fields of the Higgs doublets and singlet as
@@
$\Sigma_{h1}$ $=\frac{\delta_{1}}{48}+\frac{g_{1}^{2}}{32}+\frac{3g_{2}^{2}}{32}+\frac{\lambda_{1}}{8}+\frac{\lambda_{3}}{12}+\frac{\lambda_{4}}{24}+\frac{y_{t}^{2}}{8}+\frac{y_{b}^{2}}{8},$
$\Sigma_{h2}$ $=\frac{\delta_{2}}{48}+\frac{g_{1}^{2}}{32}+\frac{3g_{2}^{2}}{32}+\frac{\lambda_{2}}{8}+\frac{\lambda_{3}}{12}+\frac{\lambda_{4}}{24}+\frac{y_{t}^{2}}{8}+\frac{y_{b}^{2}}{8},$
@@
$\varphi_{1}=z\cos\theta\cos\beta,\quad\varphi_{2}=z\cos\theta\sin\beta,\quad\varphi_{S}=z\sin\theta+\tilde{v}_{S},$ (73)
where $\tilde{v}_{S}$ denotes the minimum on the $\varphi_{S}$ axis, which is always nonzero because of $a_{1}=0$ in our model. The HT potential (71) becomes
$V(z,\theta,\beta;T)=c_{0}+c_{1}z+\left(c_{2}+c_{2}^{\prime}T^{2}\right)z^{2}-c_{3}z^{3}+c_{4}z^{4},$ (74)
and at $T=T_{C}$, the above potential takes the form
$V\left(z,\theta,\beta;T_{C}\right)=c_{4}z^{2}\left(z-z_{C}\right)^{2},\quad z_{C}=\frac{c_{3}}{2c_{4}},$ (75)
@@
$c_{3}$ $=\frac{1}{4}\tilde{v}_{S}\sin\theta\left(\delta_{1}\cos^{2}\beta\cos^{2}\theta+\delta_{2}\cos^{2}\theta\sin^{2}\beta+d_{2}\sin^{2}\theta\right),$ (76)
@@
From this potential, $v_{C}$ and $T_{C}$ are derived, which is
$v_{C}$ $=\sqrt{\frac{{v}_{SC}^{\prime}({v}_{SC}^{\prime}-{v}_{SC})\left(\delta_{1}\cos{\beta_{C}}^{2}+\delta_{2}\sin{\beta_{C}}^{2}\right)}{\lambda_{1}\cos_{\beta_{C}}^{4}+\lambda_{2}\sin_{\beta_{C}}^{4}+2\lambda_{345}\sin_{\beta_{C}}^{2}\cos_{\beta_{C}}^{2}}}$ (78)
$T_{C}$ $=\sqrt{\frac{-m_{1}^{2}\cos^{2}{\beta}-m_{2}^{2}\sin^{2}{\beta}+2m_{3}^{2}\cos{\beta}\sin{\beta}-\frac{1}{4}{v}_{SC}^{\prime}(\delta_{1}\cos^{2}{\beta}+\delta_{2}\sin^{2}{\beta})}{\Sigma_{h1}\cos^{2}{\beta}+\Sigma_{h2}\sin^{2}{\beta}}}.$ (79)
The physical quantities evaluated by $T_{C}$ are labeled with the subscript $C$. We can find that the condition of strong first-order EWPT (67) requires large $\delta_{1}$ and $\delta_{2}$.
@@ V Numerical results
BP1 0.0069 0.0017 0.51 1.27 0.32 8.25 -4.12 -4.12 -7863 -7812 -39962 24523 110 220
BP2 0.0085 0.0020 0.53 1.27 0.32 8.25 -4.12 -4.12 -7843 -7807 -35050 34328 110 220
BP3 0.057 0.014 0.54 1.27 0.32 8.25 -4.12 -4.12 -7771 -7790 -24774 54308 110 220
Table 2: Input and output parameters.
The input and output parameters that satisfy the above constraints are summarized in Table 2. The masses of three Higgs bosons are degenerate within 1 GeV. In addition, these three benchmark points (BPs) reflect the EWPT discussion in Sec. IV. The difference between input parameters among BPs is the size of $v_{S}$. The parameter $a_{1}$ is changed to keep the value of $d_{2}$ perturbatively appropriate.
@@
The public code cosmoTransitions Wainwright (2012) is used to evaluate EWPT. In Table 3, $T_{C}$ and corresponding VEVs for BPs are shown, and strong first-order EWPT occurs in all BPs. The effect of $\lambda_{i}$ is known to be significant in the thermal loop-derived EWPT, and since the value of $\lambda_{i}$ is the same for all BPs, the Higgs VEVs and $T_{C}$ are almost identical in three cases. In addition, the small change in the singlet VEV before and after EWPT shows that the tree-level effect by the singlet introduction does not cause EWPT this time. In fact, when using the HT potential (71), EWPT becomes second-order.
@@
Next, we show the results of numerical calculations related to the DM using the public code micrOMEGAs Belanger et al. (2010, 2021). In Fig. 2, the spin-independent cross section for the scattering between the DM and nucleus $\sigma_{\mathrm{SI}}$ as a function of $m_{H_{2}}$ is shown. Here, $m_{H_{2}}$ is used once as a variable, and the other parameters are BPs in Table 2. The shaded regions are excluded by the LZ experiment. For all BPs, there is a large dip around $m_{H_{2}}\simeq 125$ GeV. This indicates that the degenerate scalar scenario is important for the suppression of DM and nucleon scattering. Furthermore, BP1 has the smallest $\delta_{1}$ and $\delta_{2}$, which describes the SU(2)L doublet-singlet mixing, and thus satisfies the experimental results outside of the degenerate region. Although the elastic neutrino-nucleon scattering makes DM detection difficult in some regions, this does not rule out the existence of the WIMP-DM O’Hare (2021).
@@
The left panel of Fig. 3 shows $\Omega_{\chi}h^{2}$ as a function of $m_{\chi}$ for all BPs in Table. 2. The value of the DM relic density $\Omega_{\mathrm{DM}}h^{2}$ must not exceed the observed value Aghanim et al. (2020):
$\Omega_{\mathrm{DM}}h^{2}=0.1200\pm 0.0012,$ (86)
which corresponds to the black dashed line in the graph. The red dots indicate the points where the calculated value equals the observed value (86).

### v2

@@ III.1 Type-I, Type-X
Therefore, in the degenerate limit of the mediator masses, the DM-quark scattering is suppressed as long as the potential (1) is used. In other words, the mixing term between $\Phi_{2}$ and $S$ is constrained by the degenerate scalar scenario. For example, suppose that the mixing terms $\frac{c_{2}}{4}\Phi_{2}^{\dagger}\Phi_{2}S$ is added to the potential (1). The couplings $C_{\chi\chi h_{2}}$ and $C_{h_{2}s}$ become
$C_{\chi\chi h_{2}}=\frac{\delta_{2}}{4}v_{2},\quad C_{h_{2}s}=\frac{\delta_{2}}{2}v_{2}v_{S}+\frac{\sqrt{2}c_{2}}{4}v_{2}.$ (67)
Then,
$A$ $=\frac{1}{2},~{}~{}~{}\Delta_{2}=-\frac{\sqrt{2}c_{2}}{4}v_{2},$ (68)
which is inconsistent with the condition (63). Thus, if the mixing terms such as $\Phi_{2}^{\dagger}\Phi_{2}S,\Phi_{2}^{\dagger}\Phi_{2}S^{2}$ are added, the suppression mechanism does not work. [^2] A more detailed analysis, including one-loop scattering, is given in Ref. Cho and Idegawa (2023). This discussion can be applied almost directly to the case of the 2HDMS.
III.2 Type-II, Type-Y
The Yukawa interaction for Type-II and Type-Y is given by
$-\mathcal{L}_{\mathrm{Yukawa}}$ $=\sum_{i=1}^{3}C_{uuH_{i}}\overline{u_{L}}u_{R}H_{i}+\sum_{i=1}^{3}C_{ddH_{i}}\overline{d_{L}}d_{R}H_{i}+\mathrm{h.c.},$ (69)
where
$C_{uuH_{i}}=\frac{m_{u}}{v_{2}}O_{2i},~{}C_{ddH_{i}}=\frac{m_{d}}{v_{1}}O_{1i}.$ (70)
The amplitude becomes
$i\mathcal{M}_{\mathrm{up}}$ $=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{u}}{v_{2}}\sum_{i=1}^{3}\frac{C_{\chi\chi H_{i}}O_{2i}}{m_{H_{i}}^{2}}$
$=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{u}}{v_{2}}\sum_{i=1}^{3}\left(C_{\chi\chi h_{1}}\frac{O_{1i}O_{2i}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{2i}}{m_{H_{i}}^{2}}\right),$ (71)
$i\mathcal{M}_{\mathrm{down}}$ $=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{d}}{v_{1}}\sum_{i=1}^{3}\frac{C_{\chi\chi H_{i}}O_{1i}}{m_{H_{i}}^{2}}$
$=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{d}}{v_{1}}\sum_{i=1}^{3}\left(C_{\chi\chi h_{1}}\frac{O_{1i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}O_{1i}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{1i}}{m_{H_{i}}^{2}}\right).$ (72)
As for case A (Type-I, Type-X), this scattering is suppressed when the three Higgs masses are equal. In order to suppress the down-type quarks scattering, it is necessary to suppress $C_{\chi\chi h_{1}}$, which can be expressed as
$C_{\chi\chi h_{1}}$ $=\frac{\delta_{1}}{4}v_{1},$ (73)
with
$\delta_{1}$ $=\frac{2}{v_{1}v_{S}}\sum_{i=1}^{3}O_{1i}O_{3i}m_{H_{i}}^{2}.$ (74)
Thus, in the degenerate scalar scenario, it is important to suppress $\delta_{2}$ for the up-type scattering and $\delta_{1}$ for the down-type scattering.
We also mention the sum rule for the degenerate scalar scenario in this case. The scattering of DM and quarks for case B (Type-II, Type-Y) is as shown in Fig. 2 and up-type quarks $u$ and down-type quarks $d$ couple to the different Higgs doublet.
Figure 2: Feynman diagrams of the DM-quark scattering in the mass and gauge eigenstates of scalar mediators. Up-type quarks $u$ couple to $\Phi_{2}$, while down-type quarks $d$ couple to $\Phi_{1}$ in case B (Type-II, Type-Y).
The scattering of up-type quarks is the same as the discussion of case A, so we focus on those of down-type quarks. As can be seen from Eq. (72), the relevant part of the suppression mechanism is
$\frac{C_{\chi\chi H_{i}}O_{1i}}{m_{H_{i}}^{2}}$ $=\frac{A}{v_{S}}\left\{(C_{h_{1}s}+\Delta_{1})O_{1i}+(C_{h_{2}s}+\Delta_{2})O_{2i}+(C_{ss}+\Delta_{3})O_{3i}\right\}\frac{O_{1i}}{m_{H_{i}}^{2}},$
$=\frac{A}{v_{S}m_{H_{i}}^{2}}\left(O_{3i}O_{1i}m_{H_{i}}^{2}+\Delta_{1}O_{1i}^{2}+\Delta_{2}O_{2i}O_{1i}+\Delta_{3}O_{3i}O_{1i}\right).$ (75)
In this case, the amplitude in the degenerate Higgs region vanishes when
$\Delta_{1}=0$ (76)
In the 2HDMS, the couplings $C_{h_{1}s}$ and $C_{\chi\chi h_{1}}$ are related to
$V_{0}$ $\supset\frac{\delta_{1}}{2}\Phi_{1}^{\dagger}\Phi_{1}|S|^{2}$
$\supset C_{\chi\chi h_{1}}h_{1}\chi^{2}+C_{h_{1}s}h_{1}s,$ (77)
where
$C_{\chi\chi h_{1}}=\frac{\delta_{1}}{4}v_{1},\quad C_{h_{1}s}=\frac{\delta_{1}}{2}v_{1}v_{S}.$ (78)
Comparing (78) with (61), we find
$A$ $=\frac{1}{2},~{}~{}~{}\Delta_{1}=0.$ (79)
Thus, in the degenerate scalar scenario, the DM-quark scattering is suppressed with the potential (1) same as case A.
III.3 Higgs search
We mention the usefulness of the degenerate scalar scenario in the Higgs search experiments at the LHC. In the 2HDMS, the couplings between the Higgs boson $H_{i}$ and the SM particle $X$ are given by the SM couplings multiplied by the mixing matrix elements $O_{1i}$. Taking the decay widths of $H_{i}$ as an example, we obtain
@@ III.3 Higgs search
where $\Gamma_{h\rightarrow XX}^{\mathrm{SM}}\left(m_{H_{i}}\right)$ is the decay width of the SM Higgs $h$ as a function of $m_{H_{i}}$. When we add up all the decay widths with $m_{H_{i}}\simeq m_{h}=125~{}\mathrm{GeV}$,
@@ IV Qualitative discussion for electroweak phase transition
We discuss the compatibility of the degenerate scalar scenario and EWPT qualitatively. As mentioned in Sec. III, the orthogonality of the mixing matrix (32) is important for the degenerate scalar scenario. Here, using the elements of the mixing matrix, the mixing parameters for SU(2)L doublet-singlet in the scalar potential can be expressed as
@@
As we discussed in Sec. III, it can be seen that the DM-quark scattering is controlled by $\delta_{1}$ and $\delta_{2}$. In other words, the degenerate scalar scenario is realized by $\delta_{1}$ and $\delta_{2}$. Strong first-order EWPT is essential for EWBG, and the condition is Arnold and McLerran (1987); Bochkarev and Shaposhnikov (1987); Funakubo and Senaha (2009);
@@
where $T_{C}$ is the critical temperature at which the effective potential has two degenerate minima, and $v_{C}=\sqrt{v_{1C}^{2}+v_{2C}^{2}}$ is the nonzero Higgs VEV at $T=T_{C}$. Possible origins of first-order EWPT are the finite-temperature boson loop and the structure of the tree-level potential. In the former case, the addition of multiple bosons is effective, and the 2HDM specializes in this type of EWPT. On the other hand, in the latter case, the existence of a single scalar is important, and the CxSM causes this type of EWPT. Since the 2HDMS has multiple bosons and a singlet scalar that does not exist in the SM, EWPT of both types can occur. First, we qualitatively analyze tree-level driven EWPT using the so-called high temperature (HT) potential, which is composed of the tree-level potential and thermal masses. In the high-temperature expansion of the finite temperature effective potential, we can obtain not only the thermal masses but also the thermal cubic terms. However, here, we only add the thermal masses to the tree-level potential since we want to check whether first-order EWPT occurs due to the effect of the doublet-singlet mixing of the tree-level. Generally, first-order EWPT would be enhanced even more by adding the thermal cubic terms.
EWPT pattern of 2HDMS is $\quantity(\langle\Phi_{1}\rangle,\langle\Phi_{2}\rangle,\langle S\rangle)=(0,0,v_{S}^{\prime})\to(v_{1},v_{2},v_{S})$, where $\langle\Phi_{i}\rangle$ and $\langle S\rangle$ are defined as the VEVs of the doublet fields $\Phi_{i}$ and the singlet field $S$, respectively. To discuss EWPT, we take the Landau gauge and assume that the charged scalar bosons do not have VEVs at any temperature in order to conserve the U(1)QED. We denote the classical background fields of the Higgs doublets and singlet as
@@
$\Sigma_{h1}$ $=\frac{\delta_{1}}{48}+\frac{g_{1}^{2}}{32}+\frac{3g_{2}^{2}}{32}+\frac{\lambda_{1}}{8}+\frac{\lambda_{3}}{12}+\frac{\lambda_{4}}{24}+\frac{y_{t}^{2}}{8}+\frac{y_{b}^{2}}{8},$ (91)
$\Sigma_{h2}$ $=\frac{\delta_{2}}{48}+\frac{g_{1}^{2}}{32}+\frac{3g_{2}^{2}}{32}+\frac{\lambda_{2}}{8}+\frac{\lambda_{3}}{12}+\frac{\lambda_{4}}{24}+\frac{y_{t}^{2}}{8}+\frac{y_{b}^{2}}{8},$ (92)
@@
$\varphi_{1}=z\cos\theta\cos\beta,\quad\varphi_{2}=z\cos\theta\sin\beta,\quad\varphi_{S}=z\sin\theta+{v}_{S}^{\prime},$ (94)
where $v_{S}^{\prime}$ denotes the minimum on the $\varphi_{S}$ axis, which is always nonzero because of $a_{1}\neq 0$ in our model. The HT potential (90) becomes
$V^{\mathrm{HT}}(z,\theta,\beta;T)=c_{0}+c_{1}z+\left(c_{2}+c_{2}^{\prime}T^{2}\right)z^{2}-c_{3}z^{3}+c_{4}z^{4}.$ (95)
Eq. (95) is composed of the tree-level potential and the thermal masses, and there is no thermal cubic term like $-c^{\prime}_{3}Tz^{3}$. This means that strong first-order EWPT could be achieved even if $c^{\prime}_{3}=0$ in this model. After shifting the vacuum energy with $c_{0}=0$, the above potential at $T=T_{C}$ takes the form
$V^{\mathrm{HT}}\left(z,\theta,\beta;T_{C}\right)=c_{4}z^{2}\left(z-z_{C}\right)^{2},\quad z_{C}=\frac{c_{3}}{2c_{4}},$ (96)
@@
$c_{3}$ $=\frac{1}{4}v_{S}^{\prime}\sin\theta\left(\delta_{1}\cos^{2}\beta\cos^{2}\theta+\delta_{2}\cos^{2}\theta\sin^{2}\beta+d_{2}\sin^{2}\theta\right),$ (97)
@@
From this potential, $v_{C}$ and $T_{C}$ are derived, which are
$v_{C}$ $=\sqrt{\frac{{v}_{SC}^{\prime}({v}_{SC}^{\prime}-{v}_{SC})\quantity(\delta_{1}\cos^{2}\beta_{C}+\delta_{2}\sin^{2}\beta_{C})}{\lambda_{1}\cos_{\beta_{C}}^{4}+\lambda_{2}\sin_{\beta_{C}}^{4}+2\lambda_{345}\sin_{\beta_{C}}^{2}\cos_{\beta_{C}}^{2}}}$ (99)
$T_{C}$ $=\sqrt{\frac{-m_{1}^{2}\cos^{2}{\beta_{C}}-m_{2}^{2}\sin^{2}{\beta_{C}}+2m_{3}^{2}\cos{\beta_{C}}\sin{\beta_{C}}-\frac{1}{4}{v}_{SC}^{\prime}(\delta_{1}\cos^{2}{\beta_{C}}+\delta_{2}\sin^{2}{\beta_{C}})}{\Sigma_{h1}\cos^{2}{\beta_{C}}+\Sigma_{h2}\sin^{2}{\beta_{C}}}}.$ (100)
The physical quantities evaluated by $T_{C}$ are labeled with the subscript $C$. We can find that the condition of strong first-order EWPT (86) requires large $\delta_{1}$ and $\delta_{2}$. If, in Eq. (100), the following relation holds
[…]
it seems that $v_{C}/T_{C}$ is independent from both $\delta_{1}$ and $\delta_{2}$ since the terms related to $\delta_{1}$ and $\delta_{2}$ are canceled out between $v_{C}$ and $T_{C}$. However, there are $\delta_{1}$ and $\delta_{2}$ in $\Sigma_{h1}$ (91) and $\Sigma_{h2}$ (92), and large $\delta_{1}$ and $\delta_{2}$ are still preferred to satisfy the condition of strong first-order EWPT.
@@ V Numerical results
BP1 0.0069 0.0017 0.51 1.27 0.32 8.25 --4.12 --4.12 --7863 --7812 --19616 4176 110 220
BP2 0.0085 0.0020 0.53 1.27 0.32 8.25 --4.12 --4.12 --7843 --7807 --12172 11450 110 220
BP3 0.057 0.014 0.54 1.27 0.32 8.25 --4.12 --4.12 --7771 --7790 195.6 29339 110 220
Table 2: Input and output parameters. These BPs reproduce the observed DM relic density (108).
The input and output parameters that satisfy the above constraints are summarized in Table 2. The masses of three Higgs bosons are degenerate within 1 GeV. In addition, these three benchmark points (BPs) reflect the EWPT discussion in Sec. IV. The difference between input parameters among BPs is the size of $v_{S}$. The parameter $a_{1}$ is changed to keep the value of $d_{2}$ perturbatively appropriate. The observed DM relic density is known as Aghanim et al. (2020)
$\Omega_{\mathrm{DM}}h^{2}=0.1200\pm 0.0012,$ (108)
and we choose the DM mass that reproduces this value at each BP.
@@
The public code cosmoTransitions Wainwright (2012) is used to evaluate EWPT. In Table 3, $T_{C}$ and corresponding VEVs for BPs are shown, and strong first-order EWPT occurs in all BPs. $v_{C}/T_{C}$ is determined roughly by the ratio of the cubic and quartic terms of the field. The bosons masses are important for the former, and $\lambda_{i}$ is important for the latter. Since the boson masses are almost the same and the value of $\lambda_{i}$ is completely the same for all BPs, the Higgs VEVs and $T_{C}$ are almost identical in three cases. In addition, the small change in the singlet VEV before and after EWPT shows that the tree-level effect by the singlet introduction does not cause EWPT this time. In fact, when using the HT potential (90), EWPT becomes second-order.
@@
Next, we show the results of numerical calculations related to the DM using the public code micrOMEGAs Belanger et al. (2010, 2021). In Fig. 3, the spin-independent cross section for the scattering between the DM and nucleus $\sigma_{\mathrm{SI}}$ as a function of $m_{H_{2}}$ is shown. Here, $m_{H_{2}}$ is used once as a variable, and the other parameters are BPs in Table 2. The shaded regions are excluded by the LZ experiment. For all BPs, there is a large dip around $m_{H_{2}}\simeq 125$ GeV. This indicates that the degenerate scalar scenario is important for the suppression of DM and nucleon scattering. Although the value of $\delta_{2}$ varies between the BPs, $a_{1}$ may be more closely controlling the overall magnitude of $\sigma_{\mathrm{SI}}$. As can be seen from Eq. (52), $\sigma_{\mathrm{SI}}$ is proportional to $a_{1}^{2}$. For example, the difference in $\sigma_{\mathrm{SI}}$ between BP1 and BP2 is roughly $10^{4}$, which is due to the fact that $a_{1}^{2}\mathrm{(BP2)}/a_{1}^{2}\mathrm{(BP1)}\simeq 10^{4}$. $a_{1}$ is important for adjusting $d_{2}$ to an appropriate value, but it also affects the magnitude of $\sigma_{\mathrm{SI}}$ in this way. On the other hand, if the Higgs masses are degenerate, the scattering is suppressed regardless of the value of $a_{1}$. In other words, the degenerate scalar scenario is also valid when $a_{1}=0$, but to avoid the domain wall problem, we should consider $a_{1}\neq 0$. Thus, Although the elastic neutrino-nucleon scattering makes DM detection difficult in some regions, this does not rule out the existence of the WIMP-DM O’Hare (2021).
@@
The left panel of Fig. 4 shows $\Omega_{\chi}h^{2}$ as a function of $m_{\chi}$ for all BPs in Table. 2. The value of the DM relic density $\Omega_{\mathrm{DM}}h^{2}$ must not exceed the observed value (108), corresponding to the black dashed line in the graph. The red dots indicate the points where the calculated value equals the observed value (108) and BPs are included.